OpenStudy (anonymous):

2x+12/3+x-12/4=13/2 X=?

hero (hero):

It's \(\frac{2x+12}{3} + \frac{x-12}{4} = \frac{13}{2}\) Correct?

OpenStudy (anonymous):

yeha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is division

hero (hero):

What happens if you muliply \(\frac{13}{2} \times \frac{2}{2}\) ?

OpenStudy (anonymous):

ammm 26/4

hero (hero):

Good, so now we have two fractions with 4 in the denominator, correct?

OpenStudy (anonymous):

response simplifies

OpenStudy (anonymous):

x=6

hero (hero):

@sami-21, don't do that again. It's rude and against CoC

hero (hero):

I'm reporting you

OpenStudy (anonymous):

hey hero rebecca made a cake today, said it in the chat

OpenStudy (anonymous):

yes tanks broO :)

OpenStudy (anonymous):

taking LCM as 12 4(2x+12)+3(x-12)=13/2*12 8x+48+3x-36=78 11x+12=78 11x=78-12 11x=66 x=66/11 x=6

hero (hero):

@sami-21, what you did was very rude. You saw that I was explaining and the you just out-right posted the answer. That's very rude and against CoC

OpenStudy (anonymous):

@hero i was answering just hit the button Post mistakenly .i am newbie and sorry for that.

hero (hero):

You mistakenly hit the post button after typing the answer. Okay.

OpenStudy (anonymous):

3x +1/2+x-1/3=15/2 x=?

hero (hero):

\(\frac{3x + 1}{2} + \frac{x-1}{3} = \frac{15}{2} \) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right

hero (hero):

So we have two fractions with the same denominator, 2. So now subtract \(\frac{3x + 1}{2} \) from both sides.

hero (hero):

You will be left with \(\frac{x-1}{3} = \frac{15}{2} - \frac{3x+1}{2}\)

hero (hero):

Next, combine the fractions with the same denominator together: \(\frac{x-1}{3} = \frac{15 - 3x - 1}{2}\)

OpenStudy (anonymous):

sorry hero and finished

hero (hero):

You finished already?

hero (hero):

You didn't let me know what you got